# “partially massless” graviton, and photon fields in de Sitter spacetime under conformal symmetry

### v1

@@ I.4 Objectives and layout
The remainder of this manuscript is structured as follows to achieve our objectives. Sect. II provides a brief introduction to the dS machinery and Dirac’s six-cone formalism. In Sect. III, we consider a symmetric rank-$2$ tensor field on the five-dimensional surface of Dirac’s six-cone in $\mathbb{R}^{6}$ and derive the possible reduced fields in dS spacetime. This approach consistently leads to the graviton (strictly massless spin-$2$), the “partially massless” graviton, and the massless vector (photon) fields in dS spacetime. Finally, we discuss our findings in Sect. IV.
@@ I.5 Convention
$\left[L^{(r)}_{ab},L^{(r)}_{cd}\right]=-\mathrm{i}\left(\eta_{ac}{L^{(r)}_{bd}}+\eta_{bd}{L^{(r)}_{ac}}-\eta_{ad}{L^{(r)}_{bc}}-\eta_{bc}{L^{(r)}_{ad}}\right)\,.$ (5)
@@ II.1.2 Rank-$1$ tensor (vector) field
Technically, in the massless case, the solution space for $Q_{1}K=0$ exhibits a singularity of the form $1/\langle Q_{1}\rangle$ (see Ref. Gazeaus1 ). To eliminate this singularity, the condition of divergencelessness on $K$ ($\partial\cdot K=0$), necessary for associating $K$ with the aforementioned massless UIR of the dS group, must be relaxed Gazeaus1 ; the field equation $Q_{1}K=0$ must be solved in an expanded solution space that includes configurations where $\partial\cdot K\neq 0$. Consequently, the equation $Q_{1}K=0$ is modified to become gauge invariant:
@@
is a solution to the field equation for any scalar field $\phi_{g}$, provided $K$ is. Then, there are three main types of solutions for $K$: gauge solutions, physical solutions that are divergenceless, and solutions that are not divergenceless. The solution space carries an indecomposable representation of the dS group, with the massless UIR $(p=1\,,\,q=1)$ as its central (physical) part (refer to Ref. Gazeaus1 for the explicit form of this indecomposable representation). By introducing a gauge-fixing parameter, denoted as $c$, the explicit form of the field equation reads as Gazeaus1 :
@@ II.1.3 Rank-$2$ tensor field
Some useful relations: Here are some useful relations necessary for verifying the aforementioned equations:
$\partial_{2}\cdot\theta\phi$ $=$ $-R^{-2}D_{1}\phi\,,$ (32)
$Q_{1}D_{1}\phi$ $=$ $D_{1}Q_{0}\phi\,,$ (33)
$\partial_{2}\cdot D_{2}D_{1}\phi$ $=$ $-\left(Q_{1}+6\right)D_{1}\phi$ (34)
$=$ $-D_{1}\left(Q_{0}+6\right)\phi\,,$
where $\phi$ is a scalar field on dS spacetime.
II.1.4 Link to intrinsic coordinates
$h_{\mu\nu}(X)=x^{a}_{\,\,,\,\mu}x^{b}_{\,\,,\,\nu}\;{\cal{K}}_{ab}\big{(}x(X)\big{)}\,,$ (35)
The following transformation applies to the covariant derivatives:
$=$ ¯$\bar{\partial}^{2}\;\phi=-R^{-2}Q_{0}\,,$
where $\phi$ denotes a dS scalar field.
@@ II.2 Dirac’s six-cone formalism
where $\Psi^{(r)}$ denotes a rank-$r$ tensor field with a specific symmetry. To further restrict the space of solutions, particularly for symmetric tensors, the above system can be supplemented, for instance, with the following CI conditions:
@@ 3. Tracelessness:
Note that: (i) The CI property of Eqs. (45), (46), (47), and (II.2) should be understood with respect to the invariance under the conformal infinitesimal transformations:
@@
where $L^{(r)}_{AB}=M_{AB}+S^{(r)}_{AB}$, with $S^{(r)}_{AB}$ acting on the indices of $\Psi^{(r)}$ in a specific permutational manner. (ii) Invariant subspaces of solutions to the CI system (45) can be readily identified by employing constant vectors $Z$, such that:
$\Psi^{(r)}_{A_{1}\ldots A_{r}}={\cal{S}}Z_{A_{1}}\ldots Z_{A_{r}}\Phi\,,\qquad\Phi\equiv\Psi^{(0)}\,,$ (50)
or gradients $\nabla$, such that:
$\Psi^{(r)}_{A_{1}\ldots A_{r}}={\cal{S}}\nabla_{A_{1}}\ldots\nabla_{A_{r-r^{\prime}}}\Psi^{(r^{\prime})}_{A_{r-r^{\prime}}\ldots A_{r}}\,,$ (51)
where, again, ${\cal{S}}$ represents appropriate symmetrizer operators.
@@ III.1 Symmetric rank-$2$ tensor field $\Psi^{(2)}_{AB}$
where $\Lambda_{b}$ and $\lambda$ represent a vector field and scalar field, respectively, both with conformal degree zero and satisfy $\left(Q_{0}-2\right)\Lambda_{b}=0$ and $\left(Q_{0}-2\right)\lambda=0$. To facilitate future analysis, let us further examine the implications of Eqs. (67) and (68):
@@
$\Rightarrow\;\;\eta^{ab}\Psi_{ab}^{(2)}+R^{-2}x\cdot\Psi^{(2)}\cdot x=\lambda-x\cdot\Lambda\,.$ (69)
Some useful relations: The following relations are essential for verifying the equations mentioned above:
$\left(Q_{0}-2\right)x_{a}=$ $\,x_{a}\left(Q_{0}-2\right)-4x_{a}-2D_{1a}\,,$ (70)
$\left(Q_{0}-2\right)D_{1a}=$ $\,D_{1a}\left(Q_{0}-2\right)+6D_{1a}$
$\,+2\left(Q_{0}+4\right)x_{a}\,,$ (71)
and for future use:
$Q_{0}\left(Q_{0}-2\right)x_{a}=$ $\,x_{a}Q_{0}\left(Q_{0}-2\right)$
$\,-4(3x_{a}+D_{1a})\left(Q_{0}-2\right)\,,$ (72)
$D_{1a}D_{1b}=$ $\,D_{1b}D_{1a}+x_{b}D_{1a}-x_{a}D_{1b}\,.$ (74)
@@ III.2 Reduction to dS spacetime
$\left(\Sigma^{2}_{2}\right)^{\prime}=$ $\,\lambda-x\cdot\Lambda\,,$ (83)
¯$\bar{\partial}\cdot\Sigma^{2}_{2}=$  $\,4R^{-2}\Sigma^{2}_{1}+\Lambda+R^{-2}x\lambda\,,$ (84)
$x\cdot\bar{\partial}\cdot\Sigma^{2}_{2}=$ $-\left(\Sigma^{2}_{2}\right)^{\prime}\,.$ (85)
From (83) and (85), in turn, a more convenient alternative form of (84) can be simply written as:
¯$\bar{\partial}\cdot\Sigma^{2}_{2}=$  $\,4R^{-2}\Sigma^{2}_{1}+xR^{-2}\left(\Sigma^{2}_{2}\right)^{\prime}\,,$ (86)
@@
It is important to emphasize that the alternative form (86) may include a term of the form $c\,\bar{\partial}\left(\Sigma^{2}_{2}\right)^{\prime}$, where $c$ is a constant, given that $x\cdot\bar{\partial}=0$. This term can be assumed to couple with the first term, $4R^{-2}\Sigma^{2}_{1}$, a fact that will be utilized in our later reasoning.
@@
We examine these dS fields in detail in the following subsections.
@@ III.3 Vector field $K\equiv\Sigma^{2}_{1}$
Through a series of straightforward computations (detailed in Appendix A), it can be shown that Eq. (III.3), subject to the condition (91), simplifies to:
@@ III.4 Rank-$2$ tensor field ${\cal{K}}\equiv\Sigma^{2}_{2}$
Note that: (i) To obtain Eq. (96), we have also used the identity $xR^{-2}{\cal{K}}^{\prime}=\Lambda+R^{-2}x\lambda$ (driven from Eqs. (84) and (86)) and the fact that $Q_{0}\left(Q_{0}-2\right)x\Lambda=0$. (ii) In the particular traceless case (where ¯$\bar{\partial}\cdot\mathcal{K}^{0}_{\cdot a}=4R^{-2}K_{a}$), multiplying Eq. (96) by $x^{a}$ yields a result that precisely matches Eq. (III.3), and subsequently, (93). (iii) Similarly, the consistency of Eq. (97) can be easily confirmed by multiplying both sides by $\eta^{ab}$, which directly reduces to Eq. (99). (iv) The above equation remains invariant under the purely conformal transformation ${\cal{K}}\mapsto{\cal{K}}+\delta{\cal{K}}$, where (see Eq. (56); noting that $\hat{N}_{5}{\cal{K}}=0$ since $\hat{N}_{5}\Psi^{(2)}=0$):
@@
where, as before, $Z_{a}$ represents the infinitesimal five-vector $Z_{a}=\varepsilon_{5a}$. (v) Regarding the latter point, it is worth noting that the dS traceless condition, ${\cal{K}}^{\prime}=0$, remains invariant under the purely conformal transformation given by Eq. (100), which ensures that the decomposition in Eq. (95) is CI.
By performing a series of straightforward calculations (outlined in Appendix B), it can be demonstrated that Eqs. (96) and (97), subject to the constraints (98) and (99), simplify to:
@@
Note that these equations should be respectively compared with Eqs. (31) and (28), which were obtained through a purely dS group-theoretical mechanism, as we will elaborate further in Sect. IV.
To conclude this section, it is important to emphasize that neither (III.4) nor (102) individually exhibit CI. However, their combination, as presented in Eqs. (96) and (97), does.
@@ IV.2 The “partially massless” graviton field
Note that the first constraint is consistent with Eq. (86), as discussed following Eqs. (86) and (87), while the remaining constraints align with the CI conditions (98) and (99). Moreover, one should notice that the constraint (107) precisely coincides with the very condition previously introduced to derive Eq. (31). By imposing these constraints, Eq. (III.4) takes the following form:
@@
Note: In a forthcoming work, we aim to conduct a comprehensive investigation of the phase-space realization of the “partially massless” graviton field, followed by an investigation of its covariant integral quantization, as has been done, for instance, in the $1+1$-dimensional anti-dS case GazOlmo . Furthermore, studying the backreaction of this field, together with the “massless” minimally coupled scalar field, on the metric within a semiclassical framework involving the dS spacetime promises to be particularly intriguing.
@@ IV.3 The graviton (strictly massless spin-$2$) field
Note: Although the above construction does not explicitly address the massless scalar field governed by the field equation $(Q_{0}-2)\phi=0$ (see Sect. II.1.1), the operator $(Q_{0}-2)$ plays a critical role that can be traced throughout the framework.
IV.4 Remarks on related literature Takook
We now review the closely related work of Takook et al. (2008) Takook . In their study, the authors used a similar framework, starting with a symmetric rank-$2$ tensor field of conformal degree zero on the cone that satisfies (61), and examined its reduction to the dS manifold. Unlike our approach, however, Takook et al. impose both the (conformal) transversality condition (46) and the (conformal) divergenceless condition (47) on the field from the outset, prior to performing the reduction.
Despite certain similarities in the initial setup, the only common result between their work and ours is the CI equation (93). This leads us to a closer examination of their computations, about which we have certain reservations. Specifically, we address concerns regarding some results reported by Takook et al., beginning with the vector field case. For clarity, we will adjust the notation used in Ref. Takook to align with ours.
In Ref. Takook , Appendix C is dedicated to deriving the reduced field equation for the dS vector field. However, Takook et al. appear to have overlooked the existence and essential role of the fourth-order differential equation (61), which governs the field on the cone and subsequently in dS spacetime (see Eq. (III.3)). Instead, they rely exclusively on the (conformal) divergenceless condition (47). While this condition is indeed compatible with (61) (and, by extension, with Eq. (III.3)) and can be considered complementary to it, it does not itself provide a basis for deriving Eq. (93).
As we demonstrated, Eq. (93) follows directly from the fourth-order differential equation (61) (or, equivalently, from Eq. (III.3)), independent of the (conformal) divergenceless condition (47). Notably, if we accept the approach of Takook et al., which relies solely on the (conformal) divergenceless condition (47), it suggests that the full degrees of freedom described by the second-order differential equation governing the photon field in dS spacetime could be encapsulated by a first-order differential equation. [Note that, for a transverse tensor field of conformal degree zero on the cone, the (conformal) divergenceless condition (47) does indeed constitute a first-order differential equation.] However, a careful examination of their computations reveals that their method effectively reduces to verifying the consistency between the (conformal) divergenceless condition (47) and Eq. (93), rather than deriving Eq. (93) itself.
In Ref. Takook , Appendix D is dedicated to deriving the reduced field equation for the dS rank-$2$ tensor field. The authors introduce the associated CI fourth-order differential equation — analogous to Eq. (96) in our paper — without detailing the decomposition process leading to Eqs. (III.4) and (102). We have reservations regarding the accuracy of their computations. Specifically, while they overlook the significance of the fourth-order differential equation (61) in Appendix C, they entirely neglect the critical role of the (conformal) divergenceless condition (47) in Appendix D, despite having initially imposed it in their calculations.
For instance, consider Eq. (D8) in Ref. Takook , which corresponds to Eq. (68) in our work. By combining Eqs. (C3) (derived directly from the (conformal) divergenceless condition (47)) and (A8) from Ref. Takook , we can reduce Eq. (D8) from:

### v2

@@ I.4 Objectives and layout
The remainder of this manuscript is structured as follows to achieve our objectives. Sect. II provides a brief introduction to the dS machinery and Dirac’s six-cone formalism. In Sect. III, we consider a symmetric rank-$2$ tensor field on the five-dimensional surface of Dirac’s six-cone in $\mathbb{R}^{6}$ and derive the possible reduced fields in dS spacetime. This approach consistently leads to the graviton (strictly massless spin-$2$), the “partially massless” graviton, and the massless vector (photon) fields in dS spacetime. Finally, we discuss our findings in Sect. IV. This paper is supplemented with four appendices, in which the mathematical details of the calculations are provided.
@@ I.5 Convention
$\left[L^{(r)}_{ab},L^{(r)}_{cd}\right]=\mathrm{i}\left(\eta_{ac}{L^{(r)}_{bd}}+\eta_{bd}{L^{(r)}_{ac}}-\eta_{ad}{L^{(r)}_{bc}}-\eta_{bc}{L^{(r)}_{ad}}\right)\,.$ (5)
@@ II.1.2 Rank-$1$ tensor (vector) field
Technically, in the massless case, the solution space for $Q_{1}K=0$ exhibits a singularity of the form $1/\langle Q_{1}\rangle$ (see Ref. Gazeaus1 ). To eliminate this singularity, the condition of divergencelessness on $K$ ($\partial\cdot K=0$), necessary for associating $K$ with the aforementioned massless UIR of the dS group, must be relaxed Gazeaus1 ; the field equation $Q_{1}K=0$ must be solved in an expanded solution space that includes configurations where $\partial\cdot K\neq 0$. As a result, the equation is replaced by the following gauge-invariant form:
@@
is a solution to the field equation for any scalar field $\phi_{g}$, provided $K$ is. Then, there are three main types of solutions for $K$: pure gauge solutions, physical solutions that are divergenceless, and non-divergenceless solutions. The full solution space forms an indecomposable representation of the dS group, with the massless UIR $(p=1\,,\,q=1)$ as its central, physical component. Notably, the space of one-particle physical states is associated with this massless UIR. For the explicit structure of this indecomposable representation, see Ref. Gazeaus1 . By introducing a gauge-fixing parameter, denoted as $c$, the explicit form of the field equation reads as Gazeaus1 :
@@ II.1.3 Rank-$2$ tensor field
Remark: In the dS QFT literature, fields are typically described using local (intrinsic) coordinates. For a detailed explanation of the relationship between these intrinsic coordinates and the ambient ones, readers are referred to Appendix B.
@@ II.2 Dirac’s six-cone formalism
where $\Psi^{(r)}$ denotes a rank-$r$ tensor field with a specific symmetry (note that $\Psi^{(0)}=\Phi$). To further restrict the space of solutions, particularly for symmetric tensors, the above system can be supplemented, for instance, with the following CI conditions:
@@ 3. Tracelessness:
Note that the CI property of Eqs. (37), (38), (39), and (II.2) should be understood with respect to the invariance under the conformal infinitesimal transformations:
@@
where $L^{(r)}_{AB}=M_{AB}+S^{(r)}_{AB}$, with $S^{(r)}_{AB}$ acting on the indices of $\Psi^{(r)}$ in a specific permutational manner.
@@ III.1 Symmetric rank-$2$ tensor field $\Psi^{(2)}_{AB}$
where $\Lambda_{b}$ and $\lambda$ represent a vector field and scalar field, respectively, both with conformal degree zero and satisfy $\left(Q_{0}-2\right)\Lambda_{b}=0$ and $\left(Q_{0}-2\right)\lambda=0$. For simplicity, we further impose on $\Lambda_{b}$ the transversality condition ($x\cdot\Lambda=0$) and the divergenceless condition (¯$\bar{\partial}\cdot\Lambda=0$). These constraints allow us to associate $\Lambda_{b}$ with the discrete UIR $(p=1\,,\,q=1)$. To aid future analysis, we now examine the consequences of Eqs. (57) and (58) in more detail:
@@
$\Rightarrow\;\;\eta^{ab}\Psi_{ab}^{(2)}+R^{-2}x\cdot\Psi^{(2)}\cdot x=\lambda\,.$ (59)
@@ III.2 Reduction to dS spacetime
$\left(\Sigma^{2}_{2}\right)^{\prime}=$ $\,\lambda\,,$ (68)
¯$\bar{\partial}\cdot\Sigma^{2}_{2}=$  $\,4R^{-2}\Sigma^{2}_{1}+\Lambda+R^{-2}x\lambda\,.$ (69)
Upon plugging $\lambda$ from Eq. (68) into Eq. (69), we obtain:
¯$\bar{\partial}\cdot\Sigma^{2}_{2}=$  $\,4R^{-2}\Sigma^{2}_{1}+\Lambda+xR^{-2}\left(\Sigma^{2}_{2}\right)^{\prime}\,,$ (70)
@@
We examine these dS fields in detail in the following subsections. For the sake of completeness, however, it is useful to briefly mention the CI equations governing the remaining fields:
  1. 1.
    Based on Eqs. (52) and (53), the vector field $\Sigma^{1}_{1}$ exhibits behavior closely analogous to that of the vector field $K\equiv\Sigma^{2}_{1}$. In particular, it satisfies the same CI equation as $K$, which will be introduced in the following discussion.
  2. 2.
    On the other hand, based on Eqs. (54)–(56), the remaining scalar fields, namely $\Sigma^{2}_{0}$, $\Sigma^{1}_{0}$, and $\Sigma^{0}_{0}$, are governed by the following common CI field equation:
[…]
@@ III.3 Vector field $K\equiv\Sigma^{2}_{1}$
Through a series of straightforward computations (detailed in Appendix C), it can be shown that Eq. (III.3), subject to the condition (78), simplifies to:
@@ III.4 Rank-$2$ tensor field ${\cal{K}}\equiv\Sigma^{2}_{2}$
Note that: (i) To obtain Eq. (83), we have also used the fact that $Q_{0}\left(Q_{0}-2\right)x\Lambda=0$ and ${\cal{K}}^{\prime}=\lambda$. (ii) In the particular traceless case (where ¯$\bar{\partial}\cdot\mathcal{K}^{0}_{\cdot a}=4R^{-2}K_{a}$; the term $\Lambda_{a}$ is assumed to couple with $K_{a}$), multiplying Eq. (83) by $x^{a}$ yields a result that precisely matches Eq. (III.3), and subsequently, (80). (iii) Similarly, the consistency of Eq. (84) can be easily confirmed by multiplying both sides by $\eta^{ab}$, which directly reduces to Eq. (86). (iv) The above equation remains invariant under the purely conformal transformation ${\cal{K}}\mapsto{\cal{K}}+\delta{\cal{K}}$, where (see Eq. (46); noting that $\hat{N}_{5}{\cal{K}}=0$ since $\hat{N}_{5}\Psi^{(2)}=0$):
@@
where, as before, $Z_{a}$ represents the infinitesimal five-vector $Z_{a}=\varepsilon_{5a}$. (v) Regarding the latter point, it is worth noting that the dS traceless condition, ${\cal{K}}^{\prime}=0$, remains invariant under the purely conformal transformation given by Eq. (87), which ensures that the decomposition in Eq. (82) is CI. (vi) Combining Eqs. (83) and (84) yields the CI equation for the full traceful field $\mathcal{K}$:
$Q_{0}\left(Q_{0}-2\right){\cal{K}}_{ab}=$ $\,\frac{1}{6}{\cal{S}}\big{[}-2D_{1a}(3x_{b}+D_{1b})\left(Q_{0}-2\right)+3(3x_{a}+D_{1a})\left(Q_{0}-2\right)(3x_{b}+D_{1b})$
$\quad\quad-7x_{a}(3x_{b}+D_{1b})\left(Q_{0}-2\right)-8(3x_{a}+D_{1a})\left(Q_{0}-2\right)x_{b}\big{]}\left(\bar{\partial}\cdot\bar{\partial}\cdot{\cal{K}}-4R^{-2}{\cal{K}}^{\prime}+\frac{1}{4}R^{-2}Q_{0}{\cal{K}}^{\prime}\right)$
$\quad\quad-\frac{1}{2}R^{-2}{\cal{S}}\left[x_{a}(3x_{b}+D_{1b})\left(Q_{0}-2\right)+(3x_{a}+D_{1a})\left(Q_{0}-2\right)x_{b}\right]{\cal{K}}^{\prime}\,.$ (88)
By performing a series of straightforward calculations (outlined in Appendix D), it can be demonstrated that Eqs. (83) and (84), subject to the constraints (85) and (86), simplify to:
@@
Note that these equations should be, respectively, compared with Eqs. (31) and (28), which were obtained through a purely dS group-theoretical mechanism, as we will elaborate further in Sect. IV.
Remark: One may observe that equations (III.4) and (90) exhibit an intriguing feature: they do not possess a smooth flat-spacetime limit. Specifically, when the dependence on the dS radius $R$ is made explicit, it becomes clear that these equations do not admit a regular transition to flat Minkowski spacetime. This behavior, however, is not unexpected, as these equations are associated with the descriptions of the “partially massless” and strictly massless graviton fields, respectively (see the following section).
[…]
It should also be noted that the “partially massless” graviton field has no Minkowskian counterpart, not even through the conformal extension described above. Therefore, the absence of a meaningful flat-spacetime limit should not come as a surprise.
Remark: It is important to emphasize that, unlike the earlier fourth-order differential equations (83) and (84), which preserve conformal invariance, the second-order differential equations (III.4) and (90) are not conformally invariant.
Remark: To conclude this section, it is also important to note that Eqs. (80) and (83) are comparable to the results presented in Ref. Takook , which were obtained through a different technical approach. This alternative method involves a more restrictive framework, imposing both divergenceless and traceless conditions. In this regard, it is worth noting that the use of these constraints on the conformal graviton field in Ref. Takook prevents the CI equations in dS spacetime from being uplifted to a manifestly six-dimensional CI form.
@@ IV.2 The “partially massless” graviton field
Note that the first constraint is consistent with Eq. (70), as it partially restricts the degrees of freedom associated with $\Lambda$. The remaining constraints are in agreement with the CI conditions (85) and (86). Moreover, one should notice that the constraint (95) precisely coincides with the very condition previously introduced to derive Eq. (31). By imposing these constraints, Eq. (III.4) takes the following form:
@@
Remark: In a forthcoming work, we aim to conduct a comprehensive investigation of the phase-space realization of the “partially massless” graviton field, followed by an investigation of its covariant integral quantization, as has been done, for instance, in the $1+1$-dimensional anti-dS case GazOlmo . Furthermore, studying the backreaction of this field, together with the “massless” minimally coupled scalar field, on the metric within a semiclassical framework involving the dS spacetime promises to be particularly intriguing.
@@ IV.3 The graviton (strictly massless spin-$2$) field
Remark: Although the above construction does not explicitly address the massless scalar field governed by the field equation $(Q_{0}-2)\phi=0$ (see Sect. II.1.1), the operator $(Q_{0}-2)$ plays a critical role that can be traced throughout the framework.
IV.4 The importance of global symmetry in dS physics